state domain
find composite functions
g(x)=3x
f(x)= 6x+9
(a)gog(x)=
(b)fof(x)=
(c)fog(x)=
(d) gof(x)=

I appreciate you assistance

Found 2 solutions by ikleyn, stanbon:
Answer by ikleyn(52795)

.
state domain
find composite functions
g(x)=3x
f(x)= 6x+9
(a)gog(x)=
(b)fof(x)=
(c)fog(x)=
(d) gof(x)=
------------------------------------------------

The domain is the set of all real numbers in all cases.

The entire number line.


Answer by stanbon(75887)

state domain
find composite functions
g(x)=3x
f(x)= 6x+9
(a)gog(x)= g[g(x)] = g(3x) = 3(3x) = 9x
Domain:: All Real Numbers
--------------------------
(b)fof(x)= f[f(x)] = f[6x+9] = 6(6x+9)+6 = 36x+60
Domain:: All Real Numbers
--------
(c)fog(x)= f[g(x)] = f[3x] = 6(3x)+9 = 18x+9
--------
(d) gof(x)= g[f(x)] = g(6x+9) = 3(6x+9) = 18x + 27
----------
Cheers,
Stan H.
-----------